# “Where am I?” Scene Retrieval with Language

### v1

@@ Abstract
Natural language interfaces to embodied AI are becoming more ubiquitous in our daily lives. This opens further opportunities for language-based interaction with embodied agents, such as a user instructing an agent to execute some task in a specific location. For example, “put the bowls back in the cupboard next to the fridge” or “meet me at the intersection under the red sign.” As such, we need methods that interface between natural language and map representations of the environment. To this end, we explore the question of whether we can use an open-set natural language query to identify a scene represented by a 3D scene graph. We define this task as “language-based scene-retrieval” and it is closely related to “coarse-localization,” but we are instead searching for a match from a collection of disjoint scenes and not necessarily a large-scale continuous map. Therefore, we present Text2SceneGraphMatcher, a “scene-retrieval” pipeline that learns joint embeddings between text descriptions and scene graphs to determine if they are matched. The code, trained models, and datasets will be made public.
@@ 1 Introduction
Localization is central to many important applications, such as, an embodied agent navigating and executing tasks in the real world. Traditional localization methods try to localize within maps such as point clouds, and large databases of images using visual sensors  [28, 29, 33, 42, 43, 45, 64, 66, 67, 69, 71, 80, 65, 81, 82, 54, 9]. Other works use signal-emitting sensors such as WiFi or GPS [39, 10, 11, 44], or sensors that track the movement of the embodied agent [17, 50, 15, 56]. Recent works [79, 76, 38] have started to explore a different modality for localization: language. This is relevant to numerous applications, e.g., semantically describing to an autonomous agent where it should go, or a human describing their location in an emergency call. All in all, localizing with language opens up further opportunities for natural interactions between humans and embodied agents.
@@
3D scene graphs have emerged as powerful and efficient representations of spaces [6], and they are grounded in language. They often describe the semantic information of a scene using natural language labels for the nodes and edges. Not only can scene graphs efficiently capture the semantics and spatial relations between objects, but they can also efficiently model the hierarchical relationship between spaces, rooms, and buildings, making them well-representative maps of the world. Furthermore, traditional image-based localization methods often require additional insight in knowing how to capture image queries that are better for the system to localize. On the other hand, with scene graphs, invoking a text-query is natural and simple for a human–one simply describes their environment. In this work, we bring scene graphs and language into the same paradigm, and propose a method to match scene graphs and open-set language descriptions of the same space. Trying to localize within scene graphs using language cues also expands existing localization literature to the natural language domain.
@@
Figure 1: Pipeline visualization. Given an open-set natural language query (left, red) and a reference map of environments represented by a set of 3D scene graphs (right, yellow), we establish text-to-scene-graph correspondences. The text and scene graph correspondences are matched according to their embeddings in a joint embedding space (blue). These embeddings are jointly learned by a joint embedding model (green). Additionally, the text-query content in brackets represent potential downstream applications for our system, such as an embodied agent performing some task.
@@
In particular, this work investigates the task of “language-based scene-retrieval.” Given a text of a scene and a set of scene graphs, the task is to identify which scene the text describes. Additionally, the text-queries are open-set natural language descriptions, meaning they do not follow a specific template or structure. We consider a scene to be a sub-portion of a larger environment such as an apartment, house, or building. We also consider scene graphs without metric information; they are “semantic scene graphs” where nodes and edges have semantic labels, and represent objects and spatial relationships respectively. Compared to point clouds or images, scene graphs naturally model the object relationships and semantics within the environment. Tasks similar to this work have recently been proposed for language-prompted retrieval of point clouds [79, 76, 38, 83], or images [47]. While these works perform fine-grained localization and output a six degrees-of-freedom (6DoF) pose, “scene-retrieval” is more in line with “coarse-localization,” which [79, 76, 38] include as an intermediary step in their pipeline. We propose a method to match open-set text-queries to 3D scene graphs, and compare results to coarse-localization methods that match based on point clouds or image data of the same scenes.
@@
In our method, dubbed Text2SceneGraphMatcher (Text2SGM), we utilize the close relation between scene graphs and language by first transforming the text-query into a graph of objects and relationships using a large language model (LLM). For the remained of this paper, we will call these “text-graphs.” We then train a joint embedding model, to jointly learn latent representations of the scene graphs and text-graphs. At inference time, we transform the text-query again into a “text-graph” and perform retrieval by calculating a matching score against all scene graphs. The scene and text-query pair with the highest score is then the match. In this work, we make the following main contributions:
@@ 2 Related Work
Localization refers to the task of determining a precise six degrees-of-freedom (DoF) pose of an object or camera within an environment. Localization methods, such as image-retrieval-based methods [28, 29, 64, 65, 67, 71, 80] and pose regression methods [37, 73, 36, 68, 54, 51] require map representations that are either a database of images [81, 82, 54, 9], point clouds [79, 76, 38], or other explicit [28, 29, 33, 42, 43, 45, 64, 66, 67, 69, 71, 80, 65] or implicit 3D representations [37, 72, 73, 36, 12, 13, 8, 18, 14, 51]. Most of these aforementioned works primarily utilize a two-step localization approach where they first coarsely localize within the large-scale environment to narrow down the search space and then perform a fine-grained localization step to output the precise 6DoF pose.
Scene Retrieval. Localizing in a large and dense map representation is typically computationally expensive and memory intensive, which is why some methods adopt a coarse-to-fine procedure [64, 25, 55, 5]. To our knowledge, the closest line of work to coarse-localization is scene retrieval. [75] proposes a cross-domain task of image based 3D scene retrieval. Given a collection of scenes and a single image, they identify the corresponding scene for that image, by calculating a graph edit distance between graph representations of the image and the 3D scene. In essence, they are coarsely localizing their query image in a map database with disjoint scenes. We perform a similar task of language based 3D scene retrieval.
@@
Scene Graphs. 3D scene representations aim to capture the geometry, semantics, and visual features of a 3D scene. Types of representations range from explicit point clouds [57] and meshes [31] to implicit surface representations such as occupancy maps [40], signed distance fields [22, 34], or neural implicit surfaces [48, 52, 53, 78]. Some 3D scene representations focus on semantic representations. Most notably, Armeni et al. [6], combine both geometry and semantics into a hierarchical scene graph representation.
Additionally, scene graphs conveniently sit at the intersection of semantics and geometry, combining scene geometry with a more natural language-based representation of the environment. With the rise of embodied AI, natural language is becoming the modality of choice to interact with these AI agents [1, 4, 60, 59, 3]. And already, scene graphs are being used for embodied AI during navigation, task execution [3, 27, 61, 35, 41], and scene manipulation [24]. Particularly in [60], scene graphs facilitate a faster semantic search over the entire scene for the objects and areas relevant to the task at hand. This shows that compared to dense point clouds and meshes, scene graphs can be lightweight while still encoding the general structure of the environment as well as the semantic information. As we hope to eventually build systems where humans can interact with these embodied AI systems, the communication channels between humans and these agents need to be grounded within a shared environment, setting the stage for semantic scene graphs. Scene graphs fulfill this need for a geometric scene representation that is also grounded with higher-level semantics from natural language.
3D Vision-Language Models. Natural language has become a prevalent intermediary between humans and intelligent agents with the rise of Large Language Models (LLMs) [16, 1] and powerful image-language embedding models [58]. Many 3D Vision-Language models have emerged for a diverse set of language-grounded tasks such as 3D object identification in a scene using natural language [19, 2, 62, 30, 26, 32]. These works aim to build models that learn mappings between referential text and objects in a scene by leveraging different types of models such as Graph Neural Networks [2, 26] or transformer-based architectures [32, 30, 62]. [26] performs node matching between a language graph and a scene graph to find specific object references. But these works operate on a scene level and query for specific objects within a scene. Other language-grounded tasks include 3D dense captioning [77, 20], visual question and answering [7], and situated reasoning [46]. All in all, these works suggest that scene representations can benefit from a connection between the underlying geometry and higher level semantic concepts, paving the way for using language to identify scenes.
@@ 3.1 Problem Formulation
Scene Retrieval. We define “language-based scene-retrieval” as the task of finding the correct scene given a natural language description of the scene. The descriptions can vary in length, but they describe subsets of the objects in a scene, as well as some spatial relationships between the objects, such as “on top of,” “next to,” “hanging from,” and “to the left of.”
We propose to tackle the scene retrieval task with a Graph Transformer network inspired by [70]. In short, we jointly train a Graph Transformer network using a contrastive loss inspired by [58] to output a vector embedding for each text description and scene graph of the same scene. To reinforce the training signal, we additionally output a matching probability between 0 and 1 as a learning signal for the loss.
Scene Graph Definition. We define a scene graph $G=(V,E)$ to be a collection of nodes $V$, where $|V|=n$, and edges $E$. The nodes are described by feature vectors $\vec{l}\in\mathbb{R}^{m}$, representing the word2vec embedding [49] of the object label, and any existing attributes. The nodes $V$ may also have associated matrix of object attributes, $\mathbf{A}\in\mathbb{R}^{t\times m}$, such as color, size, material, etc. More specifically, $\vec{l}=\text{word2vec}(l)+\frac{1}{t}\sum_{i}^{t}\text{word2vec}(\mathbf{A}_{i,:})$, which is simply the sum of the label word2vec embedding and the average attribute word2vec embedding. Edges are described by edge feature vectors $\vec{r}\in\mathbb{R}^{m}$, representing the word2vec embedding of the the spatial relationship between two objects.
@@
Figure 2: Example of Scene and Text Graphs. In the top left we see an image a living room scene. The bottom left is the corresponding semantic scene graph, where nodes represent the objects, and edges represent a spatial relationship such as “contains,” “in front of,” or “on.” The top right section shows a text description of the scene, from the human-annotated dataset. The bottom right figure shows the corresponding “text-graph” of the scene description. The left hand side represents our database of scene graphs, while the right hand side represents an incoming text-query, which we first transform into a “text-graph,” and then match with the scene graph.
@@ 3.2 Data Processing
Given the natural language sentence format of our input queries, we perform two main steps to transform the text descriptions into a standardized format before feeding it into the model. We transform the text-queries into graph representations by extracting the objects and relationships – we call these “text-graphs.” We process the scene graphs to only include edges within some distance threshold. Then for both the text-graphs and scene graphs, we vectorize the node and edge labels into fixed-length feature vectors.
ScanScribe & Human Dataset to Graphs. The first step in the pipeline involves transforming the text-queries into a “text-graph.” To create a text-graph from a description, we use a large language model (LLM), GPT-4 [1], to extract the objects, attributes, and relationships. A description may say, “There is a wooden chair next to a table,” in which case the extracted graph would have nodes representing “chair” and “table,” and an edge in between representing the relationship, “next to.” Edges are only defined if a relationship is stated in the description. We prompt the LLM to extract the relevant information from the text descriptions, and output only JSON format. Examples of the prompt, inputs, and outputs are in the supp. m. A visual example of a “text-graph” can be seen in Figure 2.
@@
Node/Edge Label Vectorization. The last step involves transforming the node and edge labels into feature vectors. For this we use word2vec [49] to simply encode each label, which is a string, into a feature vector with length 300.
@@ 3.3 Model Architecture
Figure 3: Pipeline Overview. The input to our method is a text-query and a 3D scene graph potentially matching the text. Next, we independently process these inputs to obtain graphs with word2vec embeddings as nodes for the objects. Our network then performs self- and cross-attention with a final average pooling layer to obtain separate input embeddings. These embeddings are concatenated, and a matching score is predicted by a Multi-Layer Perceptron (MLP). The joint embedding model is trained by jointly optimizing a matching loss and cosine similarity loss.
An overview of the pipeline is in Figure 3. After transforming our inputs into a standardized graph format, we pass them into the joint embedding model. The first step is an attention module with a self-attention and then a cross-attention layer. The self-attention and cross-attention layers are implemented by a Graph Transformer, with a message passing operator from [70]. In the cross-attention layer, every node from the “text-graph” is connected to every node in the scene graph, and vice versa. This ensures the message passing happens across graphs. The aggregation function after message passing is simply addition, meaning the new node feature is the sum of the results from message passing with its neighbors. The resulting graphs after the cross-attention module are average pooled across all node features, into a final graph embedding vector for each “text-graph” and scene graph, $\vec{S}_{text}$ and $\vec{S}_{scene}$. Then, $\vec{S}_{text}$ and $\vec{S}_{scene}$ are concatenated into one long embedding vector before being passed into a simple 3-layer MLP that outputs a matching probability between the two embeddings.
@@ 3.4 Contrastive Loss
where $m$ is the output matching probability, and $y_{ik}$ is the target matching probability between the $i^{th}$ scene graph and the $k^{th}$ text-graph. Our overall loss is the following:
@@
This allows us the quantify how similar are a text and a scene graph.
Table 1: The recall for matching the top $k$ out of 10 scene candidates to our text-query. These candidates are sampled from the ScanScribe (top) [83] and the collected Human (bottom) datasets. Together with the baselines, Text2Pos and CLIP2CLIP, we show three versions of our proposed method, “match-prob” and “cos-sim” employing different scoring functions, and “ret-based” obtaining matches efficiently by pre-computing the scene graph embeddings in an offline stage and running simple nearest-neighbors search. Note that our three versions are trained with the matching probability and cosine similarity in the objective function. They differ in how they calculate the matching score.

### v2

# whereami-langloc.github.io

@@ Abstract
Natural language interfaces to embodied AI are becoming more ubiquitous in our daily lives. This opens up further opportunities for language-based interaction with embodied agents, such as a user verbally instructing an agent to execute some task in a specific location. For example, “put the bowls back in the cupboard next to the fridge” or “meet me at the intersection under the red sign.” As such, we need methods that interface between natural language and map representations of the environment. To this end, we explore the question of whether we can use an open-set natural language query to identify a scene represented by a 3D scene graph. We define this task as “language-based scene-retrieval” and it is closely related to “coarse-localization,” but we are instead searching for a match from a collection of disjoint scenes and not necessarily a large-scale continuous map. We present Text2SceneGraphMatcher, a “scene-retrieval” pipeline that learns joint embeddings between text descriptions and scene graphs to determine if they are a match. The code, trained models, and datasets will be made public.
@@ 1 Introduction
3D scene graphs have emerged as powerful and efficient representations of spaces [6], and they are grounded in language. They often describe the semantic information of a scene using natural language labels for the nodes and edges. Not only can scene graphs efficiently capture the semantics and spatial relations between objects, but they can also efficiently model the hierarchical relationship between spaces, rooms, and buildings. This makes scene graphs well-representative maps of the world.
Furthermore, traditional image-based localization methods often require additional insight in knowing how to capture image queries that are better for the system to localize. On the other hand, with scene graphs, invoking a text-query is natural and simple for a human–one simply describes their environment. In this wore bring scene graphs and language into the same paradigm, and propose a method to match scene graphs and open-set language descriptions of the same space. Trying to localize within scene graphs using language cues also expands existing localization literature to the natural language domain.
Localization with scene graphs and
@@
Figure 1: Pipeline visualization. Given an open-set natural language query (left, red) and a reference map of environments represented by a set of 3D scene graphs (right, yellow), we establish text-to-scene-graph correspondences. The text and scene graph correspondences are matched according to their embeddings in a joint embedding space (blue). These embeddings are jointly learned by a joint embedding model (green). Additionally, the text-query content in brackets represent potential downstream applications for our system, they are not part of the scene description.
@@
In particular, this work investigates the task of “language-based scene-retrieval.” Given a text of a scene and a set of scene graphs, the task is to identify which scene the text describes. Additionally, the text-queries are open-set natural language descriptions, meaning they do not follow a specific template or structure. We consider a scene to be a sub-portion of a larger environment such as an apartment, house, or building. We also consider scene graphs without metric information; they are “semantic scene graphs” where nodes and edges have semantic labels, and represent objects and spatial relationships respectively. Compared to point clouds or images, scene graphs naturally model the object relationships and semantics within the environment. Tasks similar to this work have recently been proposed for language-prompted retrieval of point clouds [70, 68, 39, 74], or images [45]. While these works perform “fine-grained” localization and output a six degrees-of-freedom (6DoF) pose, “scene-retrieval” is more in line with “coarse-localization,” which [70, 68, 39] include as an intermediary step in their pipeline. We propose a method to match open-set text-queries to 3D scene graphs, and compare results to “coarse-localization” methods that match based on point clouds or image data of the same scenes.
@@
In our method, dubbed Text2SceneGraphMatcher (Text2SGM), we utilize the close relation between scene graphs and language by first transforming the text-query into a graph of objects and relationships using a large language model (LLM). For the remainder of this paper, we will call these “text-graphs.” We then train a joint embedding model, to jointly learn latent representations of the scene graphs and text-graphs. At inference time, we transform the text-query again into a “text-graph” and perform retrieval by calculating a matching score against all scene graphs. The scene and text-query pair with the highest score is then the match. In this work, we make the following main contributions:
@@ 2 Related Work
Localization refers to the task of determining a precise six degrees-of-freedom (DoF) pose of an object or camera within an environment. Localization methods, such as image-retrieval-based methods [28, 29, 56, 57, 59, 63, 71] and pose regression methods [37, 65, 36, 60, 48, 47] require map representations that are either a database of images [72, 73, 48, 10], point clouds [70, 68, 39], or other explicit [28, 29, 33, 41, 42, 43, 56, 58, 59, 61, 63, 71, 57] or implicit 3D representations [37, 64, 65, 36, 11, 12, 8, 15, 13, 47]. Most of these aforementioned works primarily utilize a two-step localization approach where they first coarsely localize within the large-scale environment to narrow down the search space and then perform a fine-grained localization step to output the precise 6DoF pose. We opt for performing “coarse-localization” because it does not require geometric information, which our scene-graphs lack by design, as we aim for a lightweight representation.
Scene Retrieval. Localizing in a large and dense map representation is typically computationally expensive and memory intensive, which is why some methods adopt a coarse-to-fine procedure [56, 22, 49, 5]. To our knowledge, the closest line of work to coarse-localization is scene retrieval. [67] proposes the task of image-based 3D scene retrieval. Given a collection of scenes and a single image, they identify the corresponding scene for that image, by calculating a graph edit distance between graph representations of the image and the 3D scene. In essence, they are coarsely localizing their query image in a map database with disjoint scenes. We perform a similar task of language-based 3D scene retrieval.
Visual Place Recognition (VPR). Another similar line of work is VPR, which aims to simply identify a place given some input query [26, 19, 27, 31, 38, 34, 25, 9]. However, image-based VPR requires high-quality images with sufficient features for matching, while text-queries work better in featureless or low-bandwidth scenarios. Images can also pose privacy concerns. Also, with language, it is easier to resolve ambiguities in place symmetry. For example, given a building with symmetric wings, one can simply describe they are in the wing to the right of the main entrance.
@@
Scene Graphs. Scene representations aim to faithfully model the geometry, semantics, and visual features of a 3D scene, and some focus on semantic representations. Most notably, Armeni et al. [6], combine both geometry and semantics into a hierarchical scene graph representation. Additionally, scene graphs conveniently sit at the intersection of semantics and geometry, combining scene geometry with a more natural language-based representation of the environment. With the rise of embodied AI, natural language is becoming the modality of choice to interact with these AI agents [1, 4, 52, 51, 3]. And already, scene graphs are being used for embodied AI during navigation, task execution [3, 24, 53, 35, 40], and scene manipulation [21]. Particularly in [52], scene graphs facilitate a faster semantic search over the entire scene for the objects and areas relevant to the task at hand. This shows that compared to dense point clouds and meshes, scene graphs can be lightweight while still encoding the general structure of the environment as well as the semantic information. As we hope to eventually build systems where humans can interact with these embodied AI systems, the communication channels between humans and these agents need to be grounded within a shared environment, setting the stage for semantic scene graphs.
3D Vision-Language Models. Natural language has become a prevalent intermediary between humans and intelligent agents with the rise of Large Language Models (LLMs) [14, 1] and powerful image-language embedding models [50]. Many 3D Vision-Language Models have emerged for a diverse set of language-grounded tasks such as 3D object identification [16, 2, 54, 30, 23, 32]. These works learn mappings between text and objects in a scene [2, 23, 32, 30, 54]. [23] performs node matching between a language graph and a scene graph to find specific object references. However, these works operate on a scene level and query for specific objects within a scene. Other language-grounded tasks include 3D dense captioning [69, 17], visual question and answering [7], and situated reasoning [44]. All in all, these works pave the way for using language to identify scenes.
@@ 3.1 Problem Formulation
Scene Retrieval. We define “language-based scene-retrieval” as the task of finding the correct scene given a natural language description of the scene. The descriptions can vary in length, but they describe subsets of the objects in a scene, as well as spatial relationships between the objects, such as “on top of,” “next to,” “hanging from,” and “to the left of.” We propose to tackle the scene retrieval task with a Graph Transformer network inspired by [62]. In short, we jointly train the network using a contrastive loss inspired by [50] to output an embedding for each text description and each scene graph input. The output embeddings have a higher cosine similarity if they are from the same scene. To reinforce the training signal, we also learn a matching probability between 0 and 1, supervised by whether or not the inputs belong to the same scene.
Scene Graph Definition. We define a scene graph $G=(V,E)$ to be a collection of nodes $V$, and edges $E$. The nodes are described by feature vectors $\vec{l}\in\mathbb{R}^{m}$, representing the word2vec embedding [46] of the object label, and any existing attributes. Each node may also have an associated matrix of object attributes, $\mathbf{A}\in\mathbb{R}^{t\times m}$, such as color, size, material, etc. More specifically, $\vec{l}=\text{word2vec}(l)+\frac{1}{t}\sum_{i}^{t}\text{word2vec}(\mathbf{A}_{i,:})$, which is simply the sum of the label word2vec embedding and the average attribute word2vec embedding. Edges are described by edge feature vectors $\vec{r}\in\mathbb{R}^{m}$, representing the word2vec embedding of the the spatial relationship between two objects.
@@
Figure 2: Example of Scene and Text Graphs. In the top left is an image of a living room scene. The bottom left is the corresponding semantic scene graph, where nodes represent the objects, and edges represent a spatial relationship such as “contains,” “in front of,” or “on.” The top right section shows a text description of the scene, from the human-annotated dataset. The bottom right figure shows the corresponding “text-graph” of the scene description. The left hand side represents our database of scene graphs, while the right hand side represents an incoming text-query, which we first transform into a “text-graph,” and then match with the scene graph.
@@ 3.2 Data Processing
We first transform the text-queries into graph representations by extracting the objects and relationships – we call these “text-graphs.” We also process the scene graphs to only include edges within some distance threshold. Then, for both the text-graphs and scene graphs, we vectorize the node and edge labels into fixed-length feature vectors.
ScanScribe & Human Dataset to Graphs. We first transform the text-queries into a “text-graph.” To create a text-graph from a description, we use a large language model (LLM), GPT-4 [1], to extract the objects, attributes, and relationships. A description may say, “There is a wooden chair next to a table,” in which case the extracted graph would have nodes representing “chair” and “table,” and an edge in between representing the relationship, “next to.” Edges are only defined if a relationship is stated in the description. We prompt the LLM to extract the relevant information from the text descriptions, and output only JSON format. Examples of the prompt, inputs, and outputs are in the supp. m. A visual example of a “text-graph” can be seen in Figure 2.
@@
Node/Edge Label Vectorization. The last step involves transforming the node and edge labels into feature vectors. For this we use word2vec [46] to simply encode each label, which is a string, into a feature vector of length 300.
@@ 3.3 Model Architecture
Figure 3: Pipeline Overview. The input to our method is a text-query and a 3D scene graph potentially matching the text. Next, we independently process these inputs to obtain graphs with word2vec embeddings as nodes for the objects. Our network then performs self- and cross-attention with a final average pooling layer to obtain embeddings. These embeddings are concatenated, and a matching probability is predicted by a Multi-Layer Perceptron (MLP). The joint embedding model is trained by jointly optimizing a matching loss and cosine similarity loss.
An overview of the pipeline is in Figure 3. After transforming our inputs into a standardized graph format, we pass them into the joint embedding model. The first step is an attention module with a self-attention and then a cross-attention layer. The self-attention and cross-attention layers are implemented with the graph transformer operator inspired by [62]. In the self-attention module, the message passing happens between edges that exist within the text- or scene-graph. In the cross-attention module, if we have a text-graph with nodes A and B connected with an edge, and a scene-graph with C and D, then during cross-attention, message passing would happen between nodes (A, C), (A, D), (B, C), (B, D). After messaging passing, the nodes are aggregated by addition. Finally, the graph nodes are average pooled across all node features, into a final graph embedding vector for each text-graph and scene graph, $\vec{S}_{text}$ and $\vec{S}_{scene}$. Then, $\vec{S}_{text}$ and $\vec{S}_{scene}$ are concatenated into one long vector before being passed into a 3-layer MLP that outputs a matching probability between the two embeddings. The overall loss is a combination of a “matching probability” loss and a “cosine similarity” loss. The same architecture is used during training and inference.
@@ 3.4 Contrastive Loss
where $m$ is the output matching probability, and $y_{ik}$ is the target matching probability between the $i^{th}$ scene graph and the $k^{th}$ text-graph. $B$ is the batch size, and $j$ indexes through the text-graphs. Our overall loss is the following:
@@
This allows us the quantify how similar are a text and scene graph.
Table 1: The recall for matching the top $k$ out of 10 scene candidates to our text-query. These candidates are sampled from the ScanScribe (top) [74] and the collected Human (bottom) datasets. Together with the baselines, Text2Pos and CLIP2CLIP, we show three versions of our pipeline, “match-prob” and “cos-sim” employing different matching score calculations, and “ret-based” (precomputing scene graph embeddings and doing retrieval-based matching). Note that the underlying joint embedding models in all three models are trained with the matching probability and cosine similarity in the loss function. They differ in how they calculate the matching score.